OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

OpenStudy (anonymous):

if y = x + 6 AND y = 2x can you see that 2x MUST be equal to x + 6 ? 2x = x+ 6 2x - x = x - 6 - x x = 6

OpenStudy (anonymous):

@eigenschmeigen, would the answer be any of these A. (6, 12) B. (–12, –6) C. (–6, –12)

OpenStudy (anonymous):

one of them

OpenStudy (anonymous):

A right?

OpenStudy (anonymous):

yes. do you understand the method?

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

no problem :)